For what values c and α is the function p, defined by p(k) = {p(k) = ck^α for k = 1, 2, 3..... and p(k)=0 otherwise a probability mass function?
Answer #1

P(k) will be probability mass function only when

\sum P(k)= 1, Therefore when \\c=\frac{1}{n} , \alpha =1 \\

and k=1,2,3....n the P(k) is a probability mass function